OpenStudy (anonymous):

The slope of a line is 1/8 . The point (-8, -5) is on the line. Determine the equation of this line.

OpenStudy (anonymous):

use the formula y-y1 = m(x-x1) where m = slope and (x1,y1) are the coordinates of the point on the line.

OpenStudy (anonymous):

i need the answer not the formula duhh

OpenStudy (anonymous):

well substitute m = 1/8 and x1 = -8 and y1=-5 in the formula! I\ll give tou a start: y - (-5) = 1/8(x - (-8))

OpenStudy (anonymous):

simplifying: y + 5 = 1/8(x+8) 8y + 40 = x+8 8y -x + 32 = 0

OpenStudy (anonymous):

can you help me learn how to plug the equations in the formula

OpenStudy (anonymous):

its a matter of substituting the correct values in the formula. m is the slope of the line and you replace m by , in this case, 1/8. The point is represented by the 2 coordinates - the first one is the horizontal coordinate reprsented by x1 and the second is the vertical coordinate y1. These 2 define the point on the line.

OpenStudy (anonymous):

ok and what about the the x and the y that is left over

OpenStudy (anonymous):

when you plug in the coordinates for x1 and y1 there is still another x and y so what do you do with those

OpenStudy (anonymous):

right - what you need to know is the x1 and x2 represent PARTICULAR VALUES of x and y - in this case -8 and -5. The x and y are VARIABLES. the equation describes the relationship between all the possible values of x and y so if you plug in a value of 2 for x you can calculate the value of y. The x1 and y1 are CONstANTS and the x and y are variables.

OpenStudy (anonymous):

You dont do anything to the x and y - the equation is the answer to the problem